7x+2y=-13
-3x-8y=-23 using the substitution method

Answer by stanbon(75887)

7x+2y=-13
-3x-8y=-23
using the substitution method
-----------
Modify:
y = -(7x+13)/2
3x + 8y = 23
------------------------
Substitute for "y" and solve for "x":
3x - 8[(7x+13)/2] = 23
--------------------
6x-8(7x+13) = 46
6x-56x-104 = 46
-50x = 150
x = -3
----
Solve for "y":
y = -(7x+13)/2
y = -(-21+13)/2
y = -(-8)/2
y = 4
=====================
Cheers,
Stan H.
================